OpenStudy (anonymous):

Write a recursive definition for the sequence 11, 8, 5, 2,… I know that the definition is \[\large a_n=a_{n-1}-3\] I just don't know how the definition works please explain

OpenStudy (fibonaccichick666):

So what is the first term?

OpenStudy (anonymous):

11 right?

OpenStudy (fibonaccichick666):

yup so let \(a_n=11\)

OpenStudy (fibonaccichick666):

sorry sorry that's in your location but the n would actually be n-1

OpenStudy (fibonaccichick666):

if you follow the sequence the a_n-1 that yields 11 would be 14

OpenStudy (fibonaccichick666):

so you would have 11=14-3

OpenStudy (fibonaccichick666):

then next term is 8=11-3

OpenStudy (fibonaccichick666):

5=8-3

OpenStudy (fibonaccichick666):

and so on and so on, do you follow?

OpenStudy (anonymous):

Ah I see now thanks!

OpenStudy (fibonaccichick666):

np have a good one!